### v1

# Slaves to the Law of Large Numbers: An Asymptotic Equipartition Property for Perplexity in Generative Language Models

Abstract
We propose a new asymptotic equipartition property for the perplexity of a large piece of text generated by a language model and present theoretical arguments for this property. Perplexity, defined as a inverse likelihood function, is widely used as a performance metric for training language models. Our main result states that the logarithmic perplexity of any large text produced by a language model must asymptotically converge to the average entropy of its token distributions. This means that language models are constrained to only produce outputs from a “typical set”, which we show, is a vanishingly small subset of all possible grammatically correct outputs. We present preliminary experimental results from an open-source language model to support our theoretical claims. This work has possible practical applications for understanding and improving “AI detection” tools and theoretical implications for the uniqueness, predictability and creative potential of generative models.
1 Introduction and Motivation.
Consider a generative model, defined as an algorithm that takes a user input ${\mathbf{X}}$ and produces an output ${\mathbf{Y}}$ that statistically resembles data from a natural source. A specific type of generative model is a large language model (LLM) whose output ${\mathbf{Y}}$ is a body of text and the input is a text user prompt ${\mathbf{X}}$. State-of-the-art LLMs (Anthropic, 2024; OpenAI, 2024) are now able to produce detailed and information-rich text outputs such as entire screenplays and book-length manuscripts from a short and simple user prompt. Furthermore, LLMs produce text that can imitate human language well-enough to pass the Turing test (Pinar Saygin et al., 2000) i.e. resembles text created by humans well enough to be convincing to human observers. In this work, we argue that an LLM is strongly constrained by statistical laws. Specifically, we show that the logarithmic perplexity of any large text produced by a language model must asymptotically converge to the average entropy of its token distributions. This means that any language model is constrained to only output text strings from a typical set, which is an exponentially vanishing subset of all possible grammatically correct strings.
1.1 Contribution: Equipartition Property for Perplexity
Perplexity, defined as an inverse likelihood function, is widely used as a performance metric for language models (Meister and Cotterell, 2021). It is closely related to the information-theoretic concepts of surprisal (Levy, 2008) and cross-entropy (Cover and Thomas, 2005b), and it also appears to capture linguistic (Miaschi et al., 2021; Gamallo et al., 2017) and cognitive (Demberg and Keller, 2008; Cohen and Pakhomov, 2020) phenomena at least partially. Many “AI detection” tools for identifying synthetic text from language models are based on observed differences between the perplexity of synthetic and natural text (Mitchell et al., 2023; Gehrmann et al., 2019).
Our main result is a generalization of the well-known Asymptotic Equipartition Theorem (AEP) (Cover and Thomas, 2005a) from information theory. The simplest version of the AEP states that a long sequence of independent and identically distributed (iid) random symbols is likely to be a typical sequence (Csiszar, 1998) defined by the property that the empirical distribution of the occurrence of different symbols within the sequence is very close to the distribution from which the symbols were drawn.
The AEP itself can be thought of as a variant of the Law of Large Numbers (LLN) (Idele, 2018) applied to a log likelihood function, and just like the LLN, there is an extensive literature on extending (Moy, 1961) the AEP to a sequence of nearly-independent but not identically distributed random variables (Nishiara and Morita, 2000; Timo et al., 2010). The outputs of non-trivial language models, however, cannot be reasonably modeled as a sequence of nearly-independent symbols (Brown et al., 2020).
2 Background and Definitions
Let M be a generative model described by ${\mathbf{Y}}=m({\mathbf{X}},{\mathbf{W}})$, where the output ${\mathbf{Y}}$ consisting of a (potentially infinite) string of tokens ${\mathbf{Y}}=[Y_{1},~{}Y_{2},~{}\dots Y_{N}~{}\dots]$ is a deterministic function of user prompt ${\mathbf{X}}$ and a pseudo-random sequence ${\mathbf{W}}$. Each token is chosen from a finite set of tokens $Y_{n}\in\mathcal{T}$. The model M can produce a number of different output strings ${\mathbf{Y}}$ for the same user prompt ${\mathbf{X}}$ corresponding to different values of the pseudo-random sequence ${\mathbf{W}}$. This defines a probability distribution $P({\mathbf{Y}}|{\mathbf{X}})$, or more formally, a sequence of probability distributions $P({\mathbf{Y}}_{N}|{\mathbf{X}})$ over ${\mathbf{Y}}_{N}\in\mathcal{T}^{N}$ where ${\mathbf{Y}}_{N}\doteq[Y_{1},~{}Y_{2},~{}\dots Y_{N}]$ is the substring of ${\mathbf{Y}}$ consisting of the first $N$ tokens.
2.1 Sequential LLM
Practical implementations of LLMs specify the probability distribution iteratively (Radford et al., 2018):
$P(Y_{1},Y_{2}|{\mathbf{X}})$ $=P(Y_{1}|{\mathbf{X}})P(Y_{2}|Y_{1},{\mathbf{X}})$ (1)
and so on. Thus, the model M first draws a random value for the first token say $Y_{1}=y_{1}$ by sampling from the distribution $P(Y_{1}|{\mathbf{X}})$. Then the model determines a distribution for the second token as a function of the initial prompt ${\mathbf{X}}$ and the randomly chosen first token $y_{1}$. Thus, the second token is randomly sampled from a distribution $P(Y_{2}|{\mathbf{X}},Y_{1}=y_{1})$ and so on. We can write
$P({\mathbf{Y}}_{N}|{\mathbf{X}})$ $=p_{1}(Y_{1})p_{2}(Y_{2})\dots p_{N}(Y_{N}),~{}\mathrm{where}~{}p_{n}(Y_{n})=P(Y_{n}|Y_{1},~{}Y_{2},~{}Y_{n-1},{\mathbf{X}})$ (2)
Given a string ${\mathbf{Y}}$, open-source LLMs can be programmed to print out the distributions $p_{n}(Y_{n})$ from which its tokens were selected. Specifically, given a user prompt ${\mathbf{X}}$ and a string of tokens ${\mathbf{Y}}_{N}\equiv[Y_{1},~{}Y_{2},~{}\dots,~{}Y_{N}]$, it is possible to get a complete listing of the distributions $p_{n}(Y_{n}),~{}n=1\dots N$. Note that complete knowledge of the $p_{n}(y)$ is not the same as complete knowledge of $P({\mathbf{Y}}|{\mathbf{X}})$ even for a fixed user prompt ${\mathbf{X}}$. As an example, the former requires knowledge of $p_{3}(y)\equiv P(y|Y_{1},Y_{2},{\mathbf{X}})$ for different values of $y\in\mathcal{T}$, but only conditioned on the specific values of $Y_{1},~{}Y_{2}$ contained in one specific string ${\mathbf{Y}}$.
Remark. Equation (2) is simply an application of the Bayes rule of probability theory and it always holds for any generative model regardless of whether the tokens $Y_{n}$ are sequentially generated. However, the conditional distributions $p_{n}(y)$ are not in general easily accessible, so while (2) is true for all generative models, it may only be useful for sequential models.
The perplexity $p_{M}({\mathbf{Y}}_{N})\doteq\prod_{n=1}^{N}p_{n}(Y_{n})^{-\frac{1}{N}}$ of a (finite length) text string ${\mathbf{Y}}_{N}=[Y_{1},~{}Y_{2},~{}\dots Y_{N}]$ for a model M is defined as the per-token inverse likelihood of the string ${\mathbf{Y}}$. It is usually more convenient to work with the log-perplexity $l_{M}({\mathbf{Y}}_{N})$:
$l_{M}({\mathbf{Y}}_{N})$ $\doteq\log_{2}\left(p_{M}({\mathbf{Y}}_{N})\right)\equiv-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(p_{n}(Y_{n}))$ (3)
2.2 A Toy Problem
Let $\alpha(y),~{}\beta(y)$ be two fixed probability distributions over the (discrete) set $\mathcal{T}$ of tokens. Consider a toy problem involving two language models A and B, that each generate a string of tokens ${\mathbf{Y}}=[Y_{1},~{}Y_{2},~{}\dots]$ where each token is generated iid from the distribution $\alpha(y)$ and $\beta(y)$ respectively. The iid assumption implies that the tokens $Y_{n}$ can be thought of as being generated by a stationary and ergodic random process [^1] Strictly speaking, a stationary random process “starts” at $n=-\infty$ rather than $n=1$; here we ignore the tokens produced by the assumed random process before $n=1$..
Consider a long string ${\mathbf{A}}_{N}\doteq[A_{1},~{}A_{2},~{}\dots,~{}A_{N}],~{}N\gg 1$ randomly generated from model A. Let $p_{\mathbf{A}}(y)$ denote the empirical distribution of $y\in\mathcal{T}$ in the string ${\mathbf{A}}_{N}$:
$p_{\mathbf{A}}(y)$ $\doteq\frac{n_{\mathbf{A}}(y)}{N}\equiv\frac{1}{N}\sum_{n=1}^{N}\mathbf{1}(A_{n}\equiv y),~{}\forall y\in\mathcal{T}$ (4)
where $\mathbf{1}(.)$ denotes the indicator function and $n_{\mathbf{A}}(y)$ is the relative frequency of token $y$ in the (long) string ${\mathbf{A}}_{N}$. The log-perplexity of string ${\mathbf{A}}_{N}$ for model A is:
$l_{A}({\mathbf{A}}_{N})$ $=-\frac{1}{N}\sum_{n=1}^{N}\log_{2}(\alpha(Y_{n}))\equiv-\sum_{y\in\mathcal{T}}\frac{n_{\mathbf{A}}(y)}{N}\log_{2}(\alpha(y))\equiv H(p_{\mathbf{A}},\alpha)$ (5)
where $H(\gamma_{1},\gamma_{2})\doteq-\sum_{y\in\mathcal{T}}\gamma_{1}(y)\log_{2}(\gamma_{2}(y))$ is the cross-entropy between two distributions $\gamma_{1}(y),~{}\gamma_{2}(y)$ over $y\in\mathcal{T}$.
It is well-known $H(\gamma_{1},\gamma_{2})\geq H(\gamma_{1})$ with equality when $\gamma_{1}\equiv\gamma_{2}$, where $H(\gamma)\doteq-\sum_{y\in\mathcal{T}}\gamma(y)\log_{2}(\gamma(y))$ is the entropy of distribution $\gamma$. Thus we have from (5):
$l_{A}({\mathbf{A}}_{N})\equiv H(p_{\mathbf{A}},\alpha)$ $\geq H(\alpha)~{}{\mathrm{with~{}equality~{}iff~{}}}p_{\mathbf{A}}(y)\equiv\alpha(y),~{}\forall y\in\mathcal{T}$ (6)
The simplest version of the classical Asymptotic Equipartition Theorem (AEP) (Cover and Thomas, 2005a) from information theory states that the log-perplexity of a long string ${\mathbf{A}}_{N}$ of iid symbols is almost always very close to the entropy $H(\alpha)$ of the distribution $\alpha(y)$ the symbols are drawn from.
Proposition 1.
Simple AEP. For a long text string ${\mathbf{A}}_{N}=[A_{1},~{}A_{2},~{}\dots,~{}A_{N}]$ of iid tokens $A_{n}$ drawn from a distribution $\alpha(y)$, the log perplexity $l_{A}({\mathbf{A}}_{N})$ as defined in (3) is close to the entropy of the distribution $H(\alpha)$ of the distribution $\alpha(y)$ with high probability:
$\lim_{N\rightarrow\infty}\Pr[|l_{A}({\mathbf{A}}_{N})-H(\alpha)|>\epsilon]$ $\equiv 0,~{}\forall\epsilon>0$ (7)
Proposition 1 can be thought of as a direct consequence of the (weak) Law of Large Numbers [^2] For all of our results in this paper, convergence is in the “weak” sense i.e. convergence in probability. applied to the random variable $l_{M}({\mathbf{A}}_{N})$:
$l_{M}({\mathbf{A}}_{N})$ $\rightarrow E\left(l_{M}({\mathbf{A}}_{N})\right)\equiv H(\alpha)$ (8)
From (5) and (6), we see that (8) itself is equivalent to $p_{\mathbf{A}}(y)\equiv\frac{n_{\mathbf{A}}(y)}{N}\rightarrow\alpha(y),~{}\forall y\in\mathcal{T}$ i.e. the empirical distribution $p_{\mathbf{A}}(y)$ for long strings ${\mathbf{A}}$ is very close to $\alpha(y)$ with high probability. Putting these observations together we have:
$l_{A}({\mathbf{A}})$ $\equiv H(p_{\mathbf{A}},\alpha)\rightarrow H(\alpha)\leq H(p_{\mathbf{A}},\beta)\equiv l_{B}({\mathbf{A}})$ (9)
Intuitively, (9) states that a long string ${\mathbf{A}}$ generated from model A is likely to have a lower perplexity for model A than for any other model B. This means that a model trained to have low perplexity over a set of reference texts will generate text strings that are statistically similar to the reference texts. This is the theoretical justification for using perplexity as a loss function for training language models and it is an excellent justification - provided that we accept the Shannon model of language i.e. the idea that languages can be reasonably modeled as a stochastic sequence of tokens (Jakobson, 1961).
2.3 A Naive Method for Detecting Text Generated by a Language Model
Given a text string ${\mathbf{Y}}$, we seek to determine whether or not it was generated by a given model M. This is the problem of “white-box” detection of AI generated text where we assume knowledge of the specific language model that potentially generated the text. A naive method for doing this is as follows: compare the log-perplexity $l_{M}({\mathbf{Y}})$ of the text ${\mathbf{Y}}$ for model M with a threshold: if $l_{M}({\mathbf{Y}})<l_{th}$, we determine that ${\mathbf{Y}}$ is likely to have been generated by model M.
Consider a long string ${\mathbf{B}}\doteq[B_{1},~{}B_{2},~{}\dots,~{}B_{N}]$ randomly generated from model B. Using the same argument as in Section 2.2, we have:
$l_{B}({\mathbf{B}})$ $\equiv H(p_{\mathbf{B}},\beta)\rightarrow H(\beta)\leq H(p_{\mathbf{B}},\alpha)\equiv l_{A}({\mathbf{B}})$ (10)
Note, however, that we have no basis to assert any relationship between $l_{A}({\mathbf{A}})$ and $l_{A}({\mathbf{B}})$. Indeed if $H(\beta)\ll H(\alpha)$ i.e. the distribution $\beta$ has significantly lower entropy than $\alpha$, it cannot be ruled out that $l_{A}({\mathbf{B}})<l_{A}({\mathbf{A}})$ i.e. a text-string generated by a different model has lower perplexity for a model than a text-string generated by itself.
The mathematical properties of cross-entropy provide a rigorous basis for using perplexity to compare the compatibility of different models with a given text string. However, this theory does not support using perplexity to compare different text strings for compatibility with a given model.
2.4 A Modest Generalization
This theoretical analysis relies on Proposition 1 which only applies to toy models that generate strings of iid tokens. Of course, the simple AEP in Proposition 1 has been extended in various ways in the literature such as the following version (see Tal et al. (2017), Appendix B.4), which we will make use of in the sequel.
Proposition 2.
Generalized AEP. Consider a language model A’ that generates text string ${\mathbf{A}}^{\prime}_{N}=[A^{\prime}_{1},~{}A^{\prime}_{2},~{}\dots,~{}A^{\prime}_{N}]$ where the tokens $A^{\prime}_{n}$ are drawn independently from a sequence of distributions $\alpha_{n}(y),~{}n=1,2,\dots$. Assuming the entropies of distributions $\alpha_{n}(y)$ are uniformly bounded i.e. $H(\alpha_{n}(y))<H_{max}<\infty,~{}\forall n$, the log perplexity $l_{A^{\prime}}({\mathbf{A}}^{\prime}_{N})$ of the string ${\mathbf{A}}^{\prime}_{N}$ as defined in (3) is close to the average entropy of the distributions $\alpha_{n}(y),~{}n=1,2,\dots,N$ with high probability:
$\lim_{N\rightarrow\infty}\Pr[|l_{A^{\prime}}({\mathbf{A}}^{\prime}_{N})-H_{A^{\prime}}(N)|>\epsilon]\equiv 0,~{}\forall\epsilon>0,~{}\mathrm{where}~{}H_{A^{\prime}}(N)\equiv\frac{1}{N}\sum_{n=1}^{N}H(\alpha_{n})$ (11)
The regularity condition that the entropies of $\alpha_{n}(y)$ are uniformly bounded is trivially true if the token dictionary $\mathcal{T}$ is a finite set: $H(p(y))\leq\log_{2}|\mathcal{T}|$ for any distribution $p(y)$ over $y\in\mathcal{T}$.
While Proposition 2 does not lend itself to an intuitive interpretation in terms of the empirical distribution of the tokens $A^{\prime}_{n}$, it too is a direct consequence of the Law of Large Numbers applied to the log-perplexity random variable. Much of the analysis in Section 2.2 can be extended to generalized models like A’ in Proposition 2 that allows each token $A^{\prime}_{n}$ to be drawn from different distributions $\alpha_{n}(y)$. However, the tokens in these models must be drawn from fixed distributions independent of past tokens. This is still quite trivial compared to modern language models where the output tokens depend in highly complex and sophisticated ways on past tokens.
3 An Equipartition Property for Perplexity
We now propose a generalization of the theory described in Section 2.2 to a more interesting class of models. Consider a language model M and a random infinitely long text string ${\mathbf{Y}}$ generated by M, whose probabilities for a given prompt ${\mathbf{X}}$ is described by (2). Specifically, given a model M and a string ${\mathbf{Y}}$, (2) defines a sequence of probability distributions $p_{n}(y)$ over the set of tokens $y\in\mathcal{T}$. We will assume that the prompt ${\mathbf{X}}$ is fixed.
Definition 1.
The empirical entropy $h_{M}({\mathbf{Y}}_{N})$ of model M for string ${\mathbf{Y}}_{N}$ is defined as:
$h_{M}({\mathbf{Y}}_{N})\doteq\frac{1}{N}\sum_{n=1}^{M}H(p_{n})$ (12)
where $H(p)\equiv-\sum_{y\in\mathcal{T}}p(y)\log_{2}(p(y))$ is the entropy of the distribution $p(y)$ over $y\in\mathcal{T}$.
Note that the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ is not the same as $\frac{1}{N}H\left(P({\mathbf{Y}}_{N}|{\mathbf{X}})\right)$; the former expression is a random variable, whereas the latter is the entropy of $P({\mathbf{Y}}_{N}|{\mathbf{X}})$ over ${\mathbf{Y}}_{N}\in\mathcal{T}^{N}$ and thus is a constant number. Indeed, we can show that
$E\left[l_{M}({\mathbf{Y}}_{N})\right]$ $\equiv E\left[h_{M}({\mathbf{Y}}_{N})\right]\equiv\frac{1}{N}H\left(P({\mathbf{Y}}_{N}|{\mathbf{X}})\right)$ (13)
where the expectation is over random text strings ${\mathbf{Y}}_{N}$.
Definition 2.
Let the log-deviation $\lambda(p)$ of a probability distribution $p(y)$ over $y\in\mathcal{T}$ be defined as:
$\lambda(p)$ $\doteq\sqrt{S(p)-(H(p))^{2}}~{}\mathrm{where}~{}S(p)\doteq\sum_{y\in\mathcal{T}}p(y)\left(\log_{2}p(y)\right)^{2}$ (14)
Note that the entropy $H(p)$ and log-deviation $\lambda(p)$ are the mean and standard deviation of the log-likelihood random variable $l_{p}(y)\equiv-\log_{2}p(y)$ under distribution $p(y)$.
Definition 3.
The log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ of a string ${\mathbf{Y}}_{N}$ for model $M$ is defined as:
$\lambda_{M}({\mathbf{Y}}_{N})$ $\doteq\frac{1}{N}\sqrt{\sum_{n=1}^{N}\lambda^{2}(p_{n})}$ (15)
We can again interpret $H(p_{n}),~{}\lambda(p_{n})$ as the conditional mean and standard deviation of $-\log_{2}(\Pr(Y_{n}))$ conditioned on ${\mathbf{Y}}_{N-1}$. Clearly, if the log-deviations $\lambda(p_{n})$ of the distributions $p_{n}(y)$ for a string ${\mathbf{Y}}$ are uniformly upper-bounded, $\lambda_{M}({\mathbf{Y}}_{N})$ asymptotically vanishes.
Preview of our main result. We want to argue that $\lambda_{M}({\mathbf{Y}}_{N})$ is the standard deviation of $l_{M}({\mathbf{Y}}_{N})-h_{M}({\mathbf{Y}}_{N})$, and therefore asymptotically vanishing $\lambda_{M}({\mathbf{Y}}_{N})$ implies $l_{M}({\mathbf{Y}}_{N})\rightarrow h_{M}({\mathbf{Y}}_{N})$. But this is very wrong - while the log-deviation $\lambda(p)$ of a fixed probability distribution $p(y)$ is a deterministic number, $\lambda_{M}({\mathbf{Y}}_{N})$ of a string ${\mathbf{Y}}_{N}$ is a random variable that takes different numerical values for different random text strings ${\mathbf{Y}}$. This considerably complicates our analysis. Fundamentally, their randomness arises from the fact that the distributions $p_{n}(y)$ are different for different strings ${\mathbf{Y}}$. This is because the tokens $Y_{n}$ in the string ${\mathbf{Y}}$ are not generated independently; instead future tokens strongly depend on earlier tokens. This is an essential property of modern language models that allows them to generate sophisticated text outputs. We now introduce the conceptual device of an auxiliary language model to get around this difficulty.
3.1 Concept of an Auxiliary Language Model
Let us consider a fixed string ${\mathbf{Y}}$. For this particular string ${\mathbf{Y}}$ and model M, we have a fixed set of distributions $p_{n}(y)$. Now define an auxiliary generative model $M^{\prime}_{\mathbf{Y}}$ that generates a random string ${\mathbf{Y}}^{\prime}\doteq[Y_{1}^{\prime},~{}Y_{2}^{\prime},~{}\dots,~{}Y_{N}^{\prime},~{}\dots]$ where the tokens $Y_{n}^{\prime}$ are generated independently from the distributions $p_{n}(y)$. This construct of the auxiliary model $M^{\prime}_{\mathbf{Y}}$ for a fixed string ${\mathbf{Y}}$ allows us to work with fixed distributions $p_{n}(y)$ which are not themselves functions of a random string.
In particular, note that the empirical entropy $h_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}^{\prime}_{N})$ and log-deviation $\lambda_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}^{\prime}_{N})$ for model $M^{\prime}_{\mathbf{Y}}$ are both deterministic functions of $N$ independent of the string ${\mathbf{Y}}^{\prime}_{N}$. Specifically, we have $h_{M^{\prime}_{\mathbf{Y}}}(N)\equiv h_{M}({\mathbf{Y}}_{N})$. Similarly, we have $\lambda_{M^{\prime}_{\mathbf{Y}}}(N)\equiv\lambda_{M}({\mathbf{Y}}_{N})$ which, since we have fixed our special string ${\mathbf{Y}}$, is now also fixed. Indeed, it is easy to show that $\lambda_{M^{\prime}_{\mathbf{Y}}}(N)$ is the standard-deviation of the random variable $l_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}^{\prime}_{N})$ the log-likelihood of the random string ${\mathbf{Y}}^{\prime}_{N}$. Then, by the Chebyshev Inequality (Cohen, 2015), we have:
$\Pr\left(\left|l_{M}({\mathbf{Y}}^{\prime}_{N})-h_{M^{\prime}_{\mathbf{Y}}}(N)\right|>\epsilon\right)$ $\leq\frac{\lambda^{2}_{M^{\prime}_{\mathbf{Y}}}(N)}{\epsilon^{2}}$ (16)
Clearly, for substrings ${\mathbf{Y}}_{N}$ of our special string ${\mathbf{Y}}$ that defines the auxiliary model $M^{\prime}_{\mathbf{Y}}$, we have by definition $l_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}_{N})\equiv l_{M}({\mathbf{Y}}_{N})$. In addition, the entropy $H(P({\mathbf{Y}}^{\prime}_{N}))$ of the random strings ${\mathbf{Y}}^{\prime}_{N}$ output by auxiliary model $M^{\prime}_{\mathbf{Y}}$ is by definition $H(P({\mathbf{Y}}^{\prime}_{N}))\equiv h_{M}({\mathbf{Y}}_{N})$.
The auxiliary model $M^{\prime}_{\mathbf{Y}}$ is much simpler than the general language model M. In particular, because the tokens $Y^{\prime}_{n}$ are generated independently of other tokens, the AEP in Proposition 2 applies to strings ${\mathbf{Y}}^{\prime}$ generated by the auxiliary model $M^{\prime}_{\mathbf{Y}}$. Indeed we note from (15) that the log-deviation $\lambda^{2}_{M^{\prime}_{\mathbf{Y}}}(N)$ vanishes for large $N$ under some regularity conditions on the distributions $p_{n}(y)$, in which case, the AEP for model $M^{\prime}_{\mathbf{Y}}$ follows from (16).
3.2 Ensemble of Auxiliary Models
We can think of a generative model M as a statistical ensemble of a large number of (simpler) auxiliary models. Each auxiliary model is capable of generating a random text string consisting of a sequence of independent tokens. We can then imagine choosing randomly from one of the auxiliary models. We now make this idea more precise.
Consider an enumeration ${\mathbf{Y}}^{(1)},~{}{\mathbf{Y}}^{(2)},~{}\dots,~{}\dots,~{}{\mathbf{Y}}^{(N|\mathcal{T}|)}$ of the possible text strings ${\mathbf{Y}}_{N}\in\mathcal{T}^{N}$. Let $M_{{\mathbf{Y}}^{(n)}}$ denote the auxiliary model associated with model M and string ${\mathbf{Y}}^{(n)}$. A sequential generative model can be represented by a probability tree where each branch in the tree represents a new token being generated. The model M as well as each of the auxiliary models $M_{{\mathbf{Y}}^{(n)}}$ can each be represented by such a probability tree. An example of a probability tree representing a generative model and a sample random string ${\mathbf{Y}}$ represented by a sequence of branches through the tree are shown in Fig. 1 which also shows the probability tree for one possible auxiliary model $M^{\prime}_{\mathbf{Y}}$.
Figure 1: Probability tree for the first 3 tokens of a language model M that selects each token $Y_{n}$ from a dictionary $\mathcal{T}$ of three tokens. Also shown in red is the probability tree of an auxiliary model for a string ${\mathbf{Y}}$ with $Y_{1}=3,~{}Y_{2}=1$ and so on.
A generative model uses a pseudo-random number generator (PRNG) that produces a sequence of random symbols ${\mathbf{W}}$ to randomly choose a sequence of branches through its probability tree. As an example, suppose ${\mathbf{W}}=[W_{1},~{}W_{2},~{}\dots W_{N}~{}\dots]$ where the $W_{n}$ are independent and uniformly distributed in $[0,1]$ and a generative model uses $W_{n}$ to generate the n’th token $Y_{n}$ according to the distribution $p_{n}(y)$ using a sampling algorithm such as the classical inversion method Martino et al. (2018). Once we fix the PRNG and the sampling method, we can in principle write down a deterministic function to describe the generative model M as ${\mathbf{Y}}_{N}=m({\mathbf{X}},{\mathbf{W}}_{N})$ where recall that ${\mathbf{X}}$ is the fixed initial prompt. Each of the auxiliary models $M_{{\mathbf{Y}}^{(n)}}$ can also similarly be described by deterministic functions ${\mathbf{Y}}_{N}=m_{{\mathbf{Y}}^{(n)}}({\mathbf{X}},{\mathbf{W}}_{N})$.
We can then define a new generative model that we call the ensemble model $M_{ensemble}$ as follows. Pick a pseudo-random sequence ${\mathbf{W}}_{1}$ and generate a string ${\mathbf{Y}}_{1}=m({\mathbf{X}},{\mathbf{W}}_{1})$ from model M. Now consider the auxiliary model $M_{{\mathbf{Y}}_{1}}$ corresponding to string ${\mathbf{Y}}_{1}$. Generate a pseudo-random string ${\mathbf{Y}}_{2}=m_{{\mathbf{Y}}_{1}}({\mathbf{X}},{\mathbf{W}}_{2})$ from model $M_{{\mathbf{Y}}_{1}}$ where ${\mathbf{W}}_{2}$ is another pseudo-random sequence. The ensemble model $M_{ensemble}$ is defined by ${\mathbf{Y}}_{2}=m_{ensemble}({\mathbf{X}},{\mathbf{W}}_{1},{\mathbf{W}}_{2})$.
The following observation is crucial to our main result: a good random number generator will almost always generate pseudo-random sequences ${\mathbf{W}}$ that belong to a typical set i.e. ${\mathbf{W}}$ that satisfies the simple AEP in Proposition 1.
Definition 4.
Typical String. An infinite text string ${\mathbf{Y}}$ (or more precisely a sequence of text strings ${\mathbf{Y}}_{N},~{}N=1,2,\dots$) is said to be typical for a model M if the log-perplexity of sufficiently long sub-strings converges to their empirical entropy for model M i.e. if $\lim_{N\rightarrow\infty}|l_{M}({\mathbf{Y}}_{N})-h_{M}({\mathbf{Y}}_{N})|=0$.
We illustrate these ideas for a simple Markov language model in Section A.2.
Proposition 3.
AEP for Perplexity. For a long text string ${\mathbf{Y}}$ i.e. $N\gg 1$ generated from a language model with a given prompt ${\mathbf{X}}$, with high probability, the log perplexity $l_{M}({\mathbf{Y}})$ is close to the empirical entropy $h_{M}({\mathbf{Y}})$. More precisely:
$\lim_{N\rightarrow\infty}\Pr[|l_{M}({\mathbf{Y}}_{N})-h_{M}({\mathbf{Y}}_{N})|>\epsilon]$ $\equiv 0,~{}\forall\epsilon>0$ (17)
Proof.
Let ${\mathbf{W}}$ denote the pseudo-random sequence that generated our string ${\mathbf{Y}}$ from model M. Thus we have ${\mathbf{Y}}=m({\mathbf{X}},{\mathbf{W}})$. With high probability, ${\mathbf{W}}$ is a typical sequence. Consider the ensemble model $M_{ensemble}$ for model M with pseudo-random sequences ${\mathbf{W}}_{1},~{}{\mathbf{W}}_{2}$ and let ${\mathbf{Y}}_{1}=m({\mathbf{X}},{\mathbf{W}}_{1})$. The classical AEP in Proposition 2 for model $M^{\prime}_{{\mathbf{Y}}_{1}}$ selected by the ensemble model requires that random strings ${\mathbf{Y}}^{\prime}_{N}$ generated by model $M^{\prime}_{{\mathbf{Y}}_{1}}$ must almost always satisfy
$\lim_{N\rightarrow\infty}|l_{M^{\prime}_{{\mathbf{Y}}_{1}}}({\mathbf{Y}}^{\prime}_{N})-h_{M^{\prime}_{{\mathbf{Y}}_{1}}}(N)|=0$ (18)
Since ${\mathbf{W}}$ is almost always a typical sequence, we can set ${\mathbf{W}}_{1}={\mathbf{W}}_{2}={\mathbf{W}}$, in which case the output of the ensemble model is exactly the same as the output ${\mathbf{Y}}$ of model M. Setting ${\mathbf{W}}_{1}\equiv{\mathbf{W}}$ we get ${\mathbf{Y}}_{1}\equiv{\mathbf{Y}}$ and
$h_{M^{\prime}_{{\mathbf{Y}}_{1}}}(N)=h_{M^{\prime}_{{\mathbf{Y}}}}(N)\equiv h_{M}({\mathbf{Y}}_{N}).$ (19)
Finally setting ${\mathbf{W}}_{2}\equiv{\mathbf{W}}$ gives ${\mathbf{Y}}^{\prime}\equiv{\mathbf{Y}}$ and
$l_{M^{\prime}_{{\mathbf{Y}}_{1}}}({\mathbf{Y}}^{\prime}_{N})=l_{M^{\prime}_{\mathbf{Y}}}({\mathbf{Y}}_{N})\equiv l_{M}({\mathbf{Y}}_{N}).$ (20)
Inserting (19, 20) into (18) proves the result. ∎
Proposition 3 asserts that with high probability a sufficiently long text string will be typical for the model that it was generated from. Note the crucial role in the above proof of the assumption that the pseudo-random sequence ${\mathbf{W}}$ is typical. We can take this construct further to establish some non-asymptotic results as well. We expect from (16) that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ will be within a few standard deviations of its mean e.g. $h_{M}({\mathbf{Y}}_{N})\pm 3\lambda_{M}({\mathbf{Y}}_{N})$ even when $N$ is not large enough for asymptotics to kick in. Finally, we will show that the set of typical strings is vanishingly small compared to all grammatically correct strings.
Proposition 4.
Let $\mathcal{T}_{M}^{N}\subset\mathcal{T}^{N}$ denote the set of substrings ${\mathbf{Y}}_{N}$ of typical strings ${\mathbf{Y}}$ for a model M. The size of the “typical set” $\mathcal{T}_{M}^{N}$ is an exponentially vanishing fraction $02^{-\gamma N},~{}\gamma>$ 0 of all grammatically correct sequences.
Proof.
If we use top-k sampling, the model M generates $k$ possible grammatically correct values for each token $Y_{n}$. The number $T_{gram}$ of grammatically correct strings is then lower-bounded as $T_{gram}>k^{N}\equiv 2^{N\log_{2}k}$. Under the (mild) assumption that for a long string ${\mathbf{Y}}_{N}$, at least some fraction of the token distributions $p_{n}(y)$ are non-uniform, we have $h_{M}({\mathbf{Y}}_{N})\leq h_{max}<\log_{2}k$. By definition, the probability of model M generating string ${\mathbf{Y}}_{N}$ is $2^{-N~{}l_{M}({\mathbf{Y}}_{N})}$, so we have for the probability $P_{typ}$ of a typical string ${\mathbf{Y}}_{N}$ generated by model M:
$P_{typ}$ $\doteq 2^{-N~{}l_{M}({\mathbf{Y}}_{N})}\rightarrow 2^{-N~{}h_{M}({\mathbf{Y}}_{N})}>2^{-N~{}h_{max}}$ (21)
The total number of typical strings cannot exceed $\left|\mathcal{T}_{M}^{N}\right|\leq 2^{N~{}h_{max}}<T_{gram}~{}2^{-N\left(\log_{2}k-h_{max}\right)}$. ∎
Proposition 4 suggests that the typical sets of two different models, each set being vanishingly small, likely have zero overlap.
4 Experiments with Open-Source Language Model
We performed a series of experiments to verify the ideas described in Section 3. The basic idea is to use a local instance of the pre-trained GPT-2 model (Radford et al., 2019) to evaluate the empirical perplexity of generating a given string using the model’s conditional probability distributions and compare the result with the log-perplexity of the string again calculated using the model’s probability distributions. Software code for the experiments is available at Bell and Mudumbai (2024).
4.1 Perplexity for GPT-2 of a String Generated by GPT-2
Our first set of experiments consisted of repeatedly generating a long text string from GPT-2 and then evaluating its log-perplexity, empirical entropy and log-deviations for the same GPT-2 model.
We initialized the prompt string and the seed of random number generator with fixed values to allow for later reproduction. We then used the top-k sampling with some fixed $k$ to generate the probability distribution for the next token. We normalized the probabilities of the top-k tokens to sum to unity. We then created and saved a table consisting of token ID, token string and selection probability of each of the top $k$ tokens. Finally, a random next token is chosen by sampling from the normalized distribution, and the new token is appended to the prompt to repeat the process for $N_{max}=512$ total tokens which is the maximum number allowed by the GPT-2 model.
The random token $Y_{n}$ chosen at each step along with the probability distribution $p_{n}(y)$ from which they were chosen were saved into a json file for later processing. We ran this experiment many times for different seed values for the random number generator, different choices of $k$ for the top-k sampling and a number of different initial prompts.
For each generated string, we can then analyze the information archived in the json file to test our claim in Proposition 3. Specifically, for each token $N\in 1\dots N_{max}$ we calculated the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and the log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ of the substring ${\mathbf{Y}}_{N}$ consisting of the first $N$ tokens of the generated string.
4.2 Perplexity for GPT-2 of an Arbitrary String
We also used a slightly modified version of the procedure described in Section 4.1 to calculate the log-perplexity and empirical entropy of an arbitrary string (that was not generated by GPT-2) on the GPT-2 model. We first choose a text string that we wish to analyze. We fix an initial fragment of the string as our prompt ${\mathbf{X}}$. We tokenize the remaining portion of the string excluding the prompt using the GPT-2 tokenizer. Let $Y_{n},~{}n=1\dots N_{max}$ denote the resulting sequence of tokens.
Stating with $n=1$, we list the probability distribution $p_{n}(y)$ of the next token $Y_{n}$ for GPT-2 using top-k sampling for the $n$’th token. We can then calculate the empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ as in Section 4.1. However, to calculate the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, we also need the probability that GPT-2 will output the actual next token $Y_{n}$ from the string ${\mathbf{Y}}$ which may not be included in the list of top-$k$ tokens [^3] Note that our choice of top-k sampling is a compromise. The number of tokens in GPT-2 is quite large ($|\mathcal{T}|\approx 40,000$). If we exhaustively list the full probability distribution of all tokens, our experiment becomes very computationally expensive and slow.. When this happens, we replaced the lowest probability entry in the list of top-k tokens with the actual next token from our string along with its probability.
The normalized probability of each of the top-k tokens along with their token-ID and token strings, as well as the probability and ID of the actual next token are saved in a json file. This json file is then processed in exactly the same way as in Section 4.1 to compute the log-perplexity $l_{M}({\mathbf{Y}}_{N})$, empirical entropy $h_{M}({\mathbf{Y}}_{N})$ and the log-deviation $\lambda_{M}({\mathbf{Y}}_{N})$ of the sub-string ${\mathbf{Y}}_{N}$ for the GPT-2 model for $N=1\dots N_{max}$.
4.3 Results and Discussion
Figure 2 shows the log-perplexity and empirical entropy for sub-strings ${\mathbf{Y}}_{N}$ for the GPT-2 model as a function of $N$ for three text strings generated by the same GPT-2 model using top-k sampling with $k=100$ for Fig. 2(a) and 2(b) and $k=4$ for Fig. 2(c). The plots also shows a range around the empirical entropy corresponding to ±$\pm\lambda_{N}({\mathbf{Y}}_{N})$ and ±$\pm 2\lambda_{N}({\mathbf{Y}}_{N})$. We see that in all three cases, the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ stays within the range $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{M}({\mathbf{Y}}_{N})$ consistently. We also see that the long-term average empirical entropy is around $3.5$ bits per token in Fig. 2(a) and Fig. 2(b), and $1.5$ bits per token in Fig. 2(c), which corresponds to approximately $2^{3.5}\approx 11$ alternative tokens to choose from for top-100 sampling and $2^{1.5}\approx 3$ tokens for top-4 sampling.
(a) GPT-2-generated string with top-100 sampling.
(b) GPT-2-generated string with top-100 sampling.
(c) GPT-2-generated string with top-4 sampling.
Figure 2: Log-perplexity and empirical entropy for GPT-2 of strings $\mathbf{Y}$ generated by the same model.
(a) Human written text.
(b) Claude 3 written text.
(c) Decreasing entropy.
Figure 3: Log-perplexity and empirical entropy for GPT-2 of selected text strings.
Figures 3(a) and 3(b) show the log-perplexity and empirical entropy for sub-strings ${\mathbf{Y}}_{N}$ for the GPT-2 model as described in Section 4.2 for a string not generated by the GPT-2 model. The string in Fig. 3(a) is a paragraph written by the authors for an early draft of this manuscript, and the string in Fig. 3(b) was generated by the Claude 3 Opus model from Anthropic (2024). While our theory does not require non-convergence of $l_{M}({\mathbf{Y}}_{N})$ to $h_{M}({\mathbf{Y}}_{N})$, we do see from Fig. 3(a) and Fig. 3(b) that the log-perplexity $l_{M}({\mathbf{Y}}_{N})$ is several standard deviations $\lambda_{M}({\mathbf{Y}}_{N})$ away from the empirical entropy $h_{M}({\mathbf{Y}}_{N})$. The contrast between Figs. 3 and 2 has obvious implications for AI text detection (Tang et al., 2024), model fingerprinting (Yu et al., 2021) and so on.
4.4 Anomalies and Limitations
While the results from our experiments described in Section 4 were always consistent with Proposition 3, they did reveal certain limitations of our setup that we now discuss.
For all three strings in Fig. 2, we see that $l_{M}({\mathbf{Y}}_{N})$ stays within the range $h_{M}({\mathbf{Y}}_{N})\pm 2\lambda_{N}({\mathbf{Y}}_{N})$, but the log-deviations $\lambda_{M}({\mathbf{Y}}_{N})$ fail to become negligibly small even at $N=512$ suggesting that the GPT-2 model is not capable of generating strings sufficiently long for asymptotics to apply.
Fig. 3(c) shows an example of an output generated by GPT-2 for which the log-perplexity and empirical entropy both monotonically decrease with string length over many tokens. This suggests that the model repeatedly produces very low entropy token distributions $p_{n}(y)$, which happens for degenerate distributions where the probability mass is concentrated in a small number of tokens. In fact, on closer examination, we find that in these cases, the model is stuck in a state where the next token distribution just picks one token deterministically. This results in output text that gets stuck in a loop, repeating the same substring until $N_{max}$ has been reached.
Fundamentally, these limitations arise from cost and computational constraints. More powerful LLMs can generate longer texts and be less likely to get stuck in bad states that produce degenerate distributions and so on. In this work, all experiments were ran on a single desktop machine (Windows 11, Intel Core i9-14900K 3.20 GHz, 96.0 GB RAM, NVIDIA GeForce RTX 3090 Ti). Additional plots and more details about these experiments are presented in Section A.1.
5 Conclusions
We proposed an asymptotic property that must be satisfied by the perplexity of any long string generated by a language model and provided theoretical and experimental arguments in support of the property. While this paper focused narrowly on an exposition of the AEP property, it opens up many questions for further study. For instance, Proposition 4 suggests that different language models have distinct and unique “model signatures.” The contrast between Fig. 2 and Fig. 3 suggests potential applications for understanding and enhancing AI detection algorithms. As reported in Section 4.4, our experimental results are limited by computational constraints and can be strengthened by improved hardware and software. A broader question is to explore other statistical laws and constraints that generative models are bound by.
Appendix A Supplemental Material
A.1 Additional Experimental Context and Results
In this work, all experiments were ran on a single desktop machine (Windows 11, Intel Core i9-14900K 3.20 GHz, 96.0 GB RAM, NVIDIA GeForce RTX 3090 Ti). On this setup, each run of a text generation (i.e., running GPT-2 to generate $N=512$ tokens, their distributions, etc.), file I/O, data analysis and plotting takes several seconds. We were able to run a few hundred of these experiments for this paper from which selected examples are presented in Section 4.
Figure 4 shows additional results similar to Fig. 2 where we naively varied the seed value of the random number generator from 100-105. We present these to show that Fig. 2’s results are representative of the general case and were not cherry-picked. In all cases, $K=100$ and $N=512$. Further, should the community like to validate or expand upon this work, we have open sourced our repository [Bell and Mudumbai, 2024]. In addition to the source code, the supplemental material includes an /examples directory that includes the target text and the anomalous generated text for the plots in Fig. 3.
(a) Seed=100.
(b) Seed=101.
(c) Seed=102.
(d) Seed=103.
(e) Seed=104.
(f) Seed=105.
Figure 4: Log-perplexity and empirical entropy for GPT-2-generated string with top-100 sampling.
A.2 A Simple Markov Language Model
We now illustrate the ideas from Section 3 for a simple Markov language model, $M_{1}$ with a binary token space consisting of only two tokens $\mathcal{T}=\{t_{1},t_{2}\}$ and a parameter $\rho\in(0,1)$. The user prompt ${\mathbf{X}}$ consists of a single token which we can set to be $t_{1}$ without loss of generality: ${\mathbf{X}}=[t_{1}]$. The model $M_{1}$ generates a text string of length $N$ following a simple probabilistic algorithm: the $n$’th token $Y_{n}$ is set equal to the $n-1$’th token with probability $\rho$. The prompt serves as the initial token i.e. $Y_{0}\doteq t_{1}$. Let $h_{b}(\rho)\doteq-\rho\log_{2}\rho-(1-\rho)\log_{2}(1-\rho)$ denote the binary entropy function with argument $\rho$.
It is easy to see that $H(p_{n})\equiv h_{b}(\rho),~{}\forall n$ and $H_{M_{1}}({\mathbf{Y}})\equiv h_{b}(\rho)$ for any text string ${\mathbf{Y}}$. For $\rho=0.9$, $H_{M_{1}}({\mathbf{Y}})\equiv h_{b}(0.9)\approx 0.47$
Example 1.
Consider the string ${\mathbf{Y}}_{1}$ with all tokens $Y_{n}=t_{1},~{}\forall n=1\dots N$. The log-perplexity is given by $l_{M_{1}}({\mathbf{Y}}_{1})=-\log_{2}(\rho)$. For $\rho=0.9$, $l_{M_{1}}({\mathbf{Y}}_{1})=-\log_{2}(0.9)\approx 0.15$.
Example 2.
Consider the string ${\mathbf{Y}}_{2}$ with alternating token values: ${\mathbf{Y}}_{2}\equiv[t_{2},t_{1},t_{2},t_{1},t_{2},t_{1}\dots]$. The log-perplexity is given by $l_{M_{1}}({\mathbf{Y}}_{2})=-\log_{2}(1-\rho)$. For $\rho=0.9$, $l_{M_{1}}({\mathbf{Y}}_{2})=-\log_{2}(0.1)\approx 3.3$.
Note that ${\mathbf{Y}}_{1}$ and ${\mathbf{Y}}_{2}$ are the most likely and least likely string of length $N$ respectively for $\rho=0.9$ (or indeed for any $\rho>0.5$). However, it is clear that neither string is typical for any reasonable $\epsilon$ with the perplexity of ${\mathbf{Y}}_{1}$ being abnormally low and that of ${\mathbf{Y}}_{2}$ abnormally high respectively.
Example 3.
Consider the string ${\mathbf{Y}}_{3}$ created by flipping a fair coin repeatedly to generate a vector whose elements are iid randomly selected with $\Pr(t_{1})=\Pr(t_{2})=0.5$. Since $Y_{3}$ is a random string, its perplexity $l_{M_{1}}({\mathbf{Y}}_{3})$ is also a random variable. However, for long strings i.e. $N\gg 1$, with high probability, $l_{M_{1}}({\mathbf{Y}}_{3})\approx-0.5\log_{2}\rho-0.5\log_{2}(1-\rho)$. For $\rho=0.9$, $l_{M_{1}}({\mathbf{Y}}_{3})\approx 1.74$ which is incidentally the arithmetic mean of $l_{M_{1}}({\mathbf{Y}}_{1})$ and $l_{M_{1}}({\mathbf{Y}}_{2})$.
Clearly, with high probability, the string ${\mathbf{Y}}_{3}$ is not typical. Indeed, from these examples, we can see that typical strings for the model $M_{1}$ are precisely those that contain close to $(1-\rho)N$ token ’flips’ i.e. instances when $Y_{n+1}\neq Y_{n}$.